Two friends part company at 11pm. The first walks north at 1.2 m/sec, while the second walks east at 1.8 m/sec. How fast is the distance between them increasing 45 seconds later? Round your answer to the nearest tenth of a m/sec.


1
Expert's answer
2021-04-14T12:47:00-0400

Let,

a= distance covered by 1st person in north in time t.

b=distance covered by 2nd person in east in time t.


Let x be the direct distance between both at time t.




And we can calculate this x using pythagoras theorem.


"\boxed {x=\sqrt{ a^2+b^2}}"

Here given,

a=1.2*45=54

b=1.8*45=81


Therefore,x=11.61


Now,



"\boxed {x^2={ a^2+b^2}}"


Now for to calculate h ow fast x is increasing, we will differentiate the function w.r.t x.


"2x{dx\over dt}=2a+2b\\"


Now at given a and b


"2x{dx\over dt}=2.88t+6.48t
\\
2x{dx\over dt}=9.36t"


Now at t= 45 and x=11.61


"{dx\over dt}={9.36\over 2*(11.61)}"


"{dx\over dt}=18.139"


After rounding off to nearest tenth,



"\boxed{{dx\over dt}=20m /s}"